### v2

# LDM: Large Tensorial SDF Model for Textured Mesh Generation

Abstract
Previous efforts have managed to generate production-ready 3D assets from text or images. However, these methods primarily employ NeRF or 3D Gaussian representations, which are not adept at producing smooth, high-quality geometries required by modern rendering pipelines. In this paper, we propose LDM, a novel feed-forward framework capable of generating high-fidelity, illumination-decoupled textured mesh from a single image or text prompts. We firstly utilize a multi-view diffusion model to generate sparse multi-view inputs from single images or text prompts, and then a transformer-based model is trained to predict a tensorial SDF field from these sparse multi-view image inputs. Finally, we employ a gradient-based mesh optimization layer to refine this model, enabling it to produce an SDF field from which high-quality textured meshes can be extracted. Extensive experiments demonstrate that our method can generate diverse, high-quality 3D mesh assets with corresponding decomposed RGB textures within seconds. The project code is available at https://github.com/rgxie/LDM.
1 Introduction
Generating 3D content is vital for diverse applications and tasks. Recent 3D generation works leverage large 2D content generation diffusion models  [24] to generate multi-view consistent images based on certain conditions, and elevate the 2D prior knowledge within diffusion models to 3d space through score distillation sampling (SDS) [22, 11] or reconstruction methods [15, 16]. Despite the impressive results, these require additional optimization time during the forward generation of 3D assets, often taking tens of minutes.
LRM  [5], on the other hand, propose an end-to-end 3D Large Reconstruction Model that predicts the 3D model of an object from a single input image in about 5 seconds, requiring no optimization time during the forward process. LRM utilizes a Transformer to encode the input image as a condition and predicts a NeRF of the object in tri-plane representation. Instant3d  [10] enhances LRM by using multi-view diffusion to generate multiple viewpoints from a single image. These views, encoded by an image encoder, serve as constraints to predict the target NeRF of the object, enabling the generation of higher-quality 3D assets from various perspectives. However, constrained by the overhead of volume rendering, LRM-based models can only be optimized under lower-resolution image patches. To achieve greater efficiency and higher resolution supervision, LGM and TMGS  [30, 40] utilize 3DGS [8] as the representation of objects, employing UNet and Transformer architectures to predict and generate the Gaussian fields of objects. However, both the NeRF and 3DGS representations struggle with issues of unsmooth geometries and suboptimal geometric quality.
SDF representation has been proven capable of reconstructing high-quality. However, introducing SDF to the learning based 3D generation pipeline encounters the challenges of high memory computation and unstable convergence. Zero12345 and Zero12345++  [13, 12] introduce 3D CNNs or 3D diffusion to predict SDF volumes from image inputs. However, the quality of the output is constrained by the resolution of the volume grid, and both 3D diffusion and 3D CNNs incur significant computational overhead. As observed in CRM  [33], predicting the correct geometry using pure RGB images is extremely hard, and it adds extra geometry information into the inputs of the pipeline to help the convergence of geometry generation with the SDF representation.
In addition to the geometry component, generating lighting-decomposed texture maps is crucial for editing and relighting 3D assets in downstream applications. However, existing learning-based 3D generation methods cannot provide lighting-decomposed texture maps, their textures embed shading effects such as shadows and highlights.
To generate 3D assets with high-quality meshes and illumination-decoupled textures within the end-to-end learning based framework, we propose LDM, a novel 3D generation pipeline with tensorial SDF representation and decoupled color field. Specifically, to improve the object surface quality and reduce memory requirements, we combine an SDF representation for the geometry with a low-rank tensorial representation in TensoRF[2]. To predict a decoupled tensorial representation, SDF values, albedo color, and shading color at any point in space are decoded from the tensorial representation and supervised by reference albedo and color images using volume rendering. To achieve finer texture and geometry details, a differentiable iso-surface extraction layer like Flexicube [26] can be integrated to enable high-resolution image supervision following the tensorial SDF representation.
However, the designed pipeline fails to converge due to facts as follows. 1) When converting SDF to density field as works like VolSDF [37], parameters like $\beta$ controlling the sharpness of conversion are required to adjust to current convergence status of SDF: a large $\beta$ results in the converted density field too smooth, causing the loss of geometric details; conversely, a small $\beta$ leads to geometric fragmentation. To address this issue, we introduce an adaptive $\beta$ adjusting schedule, making $\beta$ an optimizable parameter to allow $\beta$ to be adjusted directly by the gradient descent from the final losses. 2) Directly training our pipeline using the Flexicube render layer fails to converge due to the sparse view supervision of each object and gradients descent affecting only the vicinity of the SDF grid vertices. To enable the convergence, we leverage volume rendering which samples points in the entire space to stabilize the training and provide a initial SDF for the Flexicube render to refine local quality.
In summary, our main contributions are:
  • We propose the first feed-forward framework capable of generating high-quality triangular meshes with illumination-decoupled RGB textures from text or a single image input in just a few seconds.
  • We introduce tensorial representation into the generation of objects, accurately representing illumination-decoupled SDF fields, and enhancing convergence speed. An adaptive conversion of SDF to density strategy is proposed to enable the convergence of the novel representation in our feed-forward 3D generation framework.
  • We integrate a gradient-based mesh optimization layer to train our framework, enabling it to produce an SDF field from which high-quality triangular meshes and illumination-decoupled textures can be extracted.
Figure 1: Given a text prompt or a single image, our framework can generate corresponding high-quality 3D assets within seconds, including illumination-decoupled texture maps, facilitating integration into various downstream applications.
2 Related Work
Diffusion Models for Multi-view Synthesis
Zero123 [15] demonstrates that large diffusion models have learned rich 3D priors about the visual world, even though they are only trained on 2D images. Furthermore, they introduce the first refined diffusion model capable of generating plausible images of objects from any viewpoint. A series of subsequent works [27, 28, 16, 17] have focused on enhancing the 3D consistency and resolution of multi-view images generated by diffusion models. Zero123++  [27] proposes that generating consistent multi-view images hinges on accurately modeling their joint distribution and integrates six images in a 3×2 layout into a single frame for simultaneous multi-view generation during one diffusion process. MVDream [28] and Wonder3D [17] both recommend implementing multi-view attentions to enable feature sharing during the multi-view generation process. Free3D [39] introduces Plücker Embeddings of pixels in images’ corresponding rays and a Ray Conditional Normalization layer during the diffusion generation process, which helps to produce multi-view images with more accurate perspectives. Although these methods have generated impressive multi-view images, achieving 3D consistency among them remains a challenge.
Lifting 2D Diffusion for 3D Generation
Recent breakthroughs in multi-view diffusion model  [24] have rapidly sparked interest  [22, 15] in reconstructing 3D content generation from various conditions, such as Text and images, known as 2D-lifting, include distillation-based approaches and reconstructed-based approaches. Distillation-based approaches, exemplified by Dreamfusion and Magic3D  [22, 11], leverage 2D diffusion models for optimizing 3D representations through score distillation sampling (SDS). The reconstructed-based approaches are notably represented by Zero123 and SyncDreamer [15, 16], which involves generating images with multi-view consistency using diffusion models, followed by employing volume rendering techniques to optimize 3D representations, such as neural radiance fields (NeRF)  [19]or signed distance functions (SDF) [31]. These methods have achieved impressive results, but there are still some issues. The distillation approach tends to struggle with multi-faceted problems and requires a longer time to refine a single shape. While Zero123 [15] is capable of producing images from any viewpoint using a single input image and its relative pose, it faces challenges in maintaining consistency across multi-view images, leading to 3D representations with blurred geometric details. Moreover, these methods require additional optimization time during the forward generation of 3D assets.
Feed-forward 3D Generative Models
Currently, there are some feed-forward methods capable of generating 3D assets directly from text and image inputs without optimization. LRM [5] introduced the first end-to-end 3D Large Reconstruction Model, predicting a NeRF of the object from a single image in just 5 seconds. Since LRM only uses a single image as the condition for the transformer, the resulting object’s back side exhibits sparse coloring, inconsistent with the front. Instant3D [10] improves LRM by generating multiple viewpoints from a single image through multi-view diffusion. The encoded views act as constraints to predict the object’s NeRF, enhancing the quality of 3D assets from different perspectives. 3DGS [8] has emerged as a new attempt at object representation because it can quickly produce high-quality, high-resolution images from any viewpoint through splatting and rasterization. LGM [30]encodes attributes of 3D Gaussians from multiple views into splatter images [29] and uses a UNet to predict these images from RGB inputs.TMGS [40] uses a transformer to predict tri-plane attribute fields, enriching point cloud models of objects with additional properties to achieve a 3DGS representation. Nevertheless, both 3DGS and NeRF representations encounter difficulties with issues of uneven geometries and suboptimal geometric quality. SDF representation has been proven capable of reconstructing high-quality geometries Zero12345 and Zero12345++  [13, 12] use 3D CNNs or 3D diffusion to predict SDF volumes from images, but the output quality is limited by the volume grid’s resolution. CRM [33] uses diffusion to predict six Canonical Coordinates Maps from a single input and optimizes a UNet to predict a tri-plane from these CCMS, which includes an SDF field. CRM integrates a differentiable iso-surface extraction layer [26] to optimize the generated SDF. These efforts have produced impressive textured meshes; however, their textures embed shading effects such as shadows and highlights, making these meshes unsuitable for relighting and material editing. This limitation hinders their use in downstream applications.
3 Method
Figure 2: The overview of our framework. When given an image or text prompt condition, we first utilize a diffusion model to generate multiple viewpoint images. These images are then encoded into image feature tokens using the DINO2 image encoder. Subsequently, these tokens are fed into a transform-based tensorial object reconstructor, resulting in a tensorial SDF representation. The tensorial SDF representation can be further rendered using volume rendering or the Flexicube render layer to produce images or extract meshes.
We propose the LDM, a framework as shown in Figure 2 that takes a single image or a text prompt as input and generates a corresponding triangular mesh equipped with illumination-decoupled textures. We first employ a multi-view diffusion model to generate multiple viewpoint images of the target object, conditioned on either an input text prompt or image, as detailed in Section 3.1. In Section 3.2, we discuss the use of tensorial SDF representation for generating objects, which offers robust expressive power. Next, we introduce our two-stage training strategy. In Section 3.3, we introduce how to train a transformer-based model to generate tensorial SDF representations from sparse multiple views using volume rendering in the first stage. Finally, in the second stage detailed in  Section 3.4, we introduce a gradient-based mesh optimization layer, Flexicube [26], to refine the generative model and enhance the quality of the extracted textures.
3.1 Conditional Multi-view Generation
Since Zero123 [15], numerous diffusion models capable of producing conditional multi-view images have emerged, such as MVdream, Imagedream, SyncDreamer, and Zero123Plus [28, 32, 16, 27]. We utilize MVDream for text inputs and ImageDream for image inputs. Both models are engineered to generate multi-view images from four orthogonal azimuths at a fixed elevation, which will serve as input conditions for the subsequent generation pipeline. Compared to other methods, MVDream and ImageDream employ joint training with synthetic data from the Objaverse dataset [3] and real data from the large-scale text-to-image (t2i) dataset, LAION5B [25]. Thus, We choose these models in our framework, aiming to boost our generalization capabilities for realistic object generation.
Although the implementation of a cross-view attention mechanism in those models has improved the consistency of the generated multi-view images, yielding impressive results, subtle 3D inconsistencies may still occur. To simulate this perturbation, inspired by previous methods [30, 5], we randomly apply a grid distortion to the input views and introduce a camera jitter to the camera poses during training. By deliberately introducing these inconsistencies during training, our framework is better equipped to handle the 3D inconsistencies encountered in images generated through multi-view diffusion.
3.2 Tensorial SDF and Decoupled Color Field
Recently, many studies have proposed various implicit neural representations for 3d objects or scenes, combined with differentiable rendering techniques, achieving impressive reconstruction results. NeRF [19] employs an MLP to encode the radiance field of a scene object and uses volume rendering to transform these volume densities and colors within the radiance field into an image from the target viewpoint. Unlike NeRF, TensoRF [2] represents the scene’s radiance field with a 4D tensor. This tensor describes a 3D voxel grid, each voxel containing multi-channel features. TensoRF further decomposes this 4D tensor into several compact, low-rank components, greatly reducing memory requirements and improving rendering quality. NeRF-based methods often suffer from the issue of uneven geometries in reconstructed object surfaces. To address this, works such as VolSDF and NeuS [35, 31] propose replacing density with SDF in implicit representations to enhance the geometric quality of the reconstructions, resulting in smoother surfaces. Furthermore, TensoSDF [9], which combines tensorial representation with SDF, has proven effective in encoding various material features, such as albedo and metallic properties, in inverse rendering tasks. Inspired by these studies, we propose employing a tensorial SDF representation to depict both the illumination-decoupled color field and SDF field of the objects generated in this work.
Tensorial representation. We utilize the Vector-Matrix factorization method proposed by TensoRF as our tensorial representation. Specifically, our tensorial representation is formulated as follows:
$V_{p}=V^{X}_{k}\circ M^{YZ}_{k}\oplus V^{Y}_{k}\circ M^{ZX}_{k}\oplus V^{Z}_{k}\circ M^{XY}_{k}$ (1)
where $V_{p}$ denoted the feature vector of position $p$. In addition, $V^{m}_{k}$ and ${M}_{k}^{\tilde{m}}$ represent the $k$-th vector and matrix factors of their corresponding spatial axes $m$, and $\tilde{m}$ is the two axes orthogonal to m (e.g., $\tilde{X}=YZ$). $\circ$ and $\oplus$ denote the element-wise multiplication and concatenation operations. Moreover, unlike TensoRF, which uses two separate tensor fields to encode geometry and appearance, we leverage a single shared tensorial field for both. This approach enhances the correlation between geometry and appearance.
Model SDF and decoupled color field. We aim to decompose the final rendering color into two parts. The first part is the albedo color, which reflects the inherent color of the object and remains constant for each object. The second part is the shading color, which is determined by the varying lighting conditions the object receives and its different material properties. Given the continuous nature of the decomposed appearance field and SDF field, we introduce multiple MLP decoders to decode the purely explicit features extracted from the tensorial representation. This approach not only prevents the unsmoothness that might arise from purely explicit representations but also allows for the prediction of various material properties and SDF from compact features. Formally,
$s=\Theta_{s}(V_{p},p),(c_{a},c_{r})=\Theta_{c}(V_{p},p),$ (2)
where $\Theta_{s},\Theta_{c}$ where represent the MLPs used to decode the SDF value, albedo color, and shading color. Equipped with the albedo color $c_{a}$ and shading colors $c_{s}$, the final rendering color $c$ of point $p$ can be calculated as $c=c_{a}\cdot c_{r}$. Once the albedo map with decoupled illumination is obtained, the object can be used in various downstream applications. In new scenes, the shading color can be calculated based on different materials and lighting conditions, following the rendering equation [7].
Adaptive conversion of SDF to density. Currently, we can obtain the SDF of any point on the object, but SDF cannot be directly used for volume rendering. Follow VolSDF [35], we model the conversion process from SDF value $s$ to the corresponding density $\sigma$ by the following formula:
$\sigma=\begin{dcases}\frac{1}{\beta}\Bigl{(}1-\frac{1}{2}exp\Bigl{(}\frac{s}{\beta}\Bigr{)}\Bigr{)}&\text{if}\ s<0,\\ \frac{1}{2\beta}exp\Bigl{(}-\frac{s}{\beta}\Bigr{)}\Bigr{)}&\text{if}\ s\geq 0,\\ \end{dcases}$ (3)
where $\beta>0$ is an important parameter that controls the sharpness of the conversion from SDF to density. A large beta will result in the density field obtained from the SDF conversion being too smooth, causing the loss of geometric details. Conversely, a small beta will lead to geometric fragmentation.
Figure 3: Comparing model training performance across different Beta schedules.
To determine the appropriate beta value, several schemes are conducted. The first scheme is using the empirical mean of beta values from the reconstructed models. We reconstructed 200 scenes from multi-view images using VolSDF [37] and collected the optimized beta values from the 200 VolSDF models. The mean of these values sets the beta value to train our network. However, starting with a too-small beta made the network ultimately fail to converge. The second scheme is a linear scheduling mechanism for beta to increase training stability, following the approach of BakeSDF [36]. We initialize beta with a fixed value and gradually decrease it as the training epochs progressed, as $\beta=\beta_{0}\left(1+\frac{\beta_{0}-\beta_{1}}{\beta_{1}}t^{-1}\right)$, where $t$ goes from 0 to 1 during training, $\beta_{0}=0.1$ , and $\beta_{1}$ for the three ray points hierarchical sampling is 0.015, 0.003, and 0.001, respectively. Unfortunately, this also failed. We attribute it to that the manually set beta decay strategy might not align with the convergence rate of our generation network. Therefore, we introduce an adaptive beta adjust schedule and allow the beta value to be adjusted directly by the gradient descent process instead of a manual schedule. Specifically, we initially set beta to a relatively large value of 0.1 and made it an optimizable parameter. We train it together with the image encoder and tensorial object reconstructor within the framework. We found that this approach allows beta to gradually decrease as the model acquired more knowledge, eventually converging with the model. As seen in Figure 3, this technique effectively guides the model toward convergence.
3.3 Feed-forward Large Reconstruction Model
In this section, we describe how a transformer-based model reconstructs a tensorial representation from sparse view images generated by multi-view diffusion. Reconstructing 3D representations from sparse view inputs has always been a challenge in the community, with traditional methods struggling to incorporate and apply prior 3D knowledge. In recent years, diffusion models trained on vast datasets have demonstrated a robust understanding of prior knowledge, capable of generating images under various controlled conditions. This has sparked interest in exploring large models for 3D generation. Inspired by Instant3D and LRM  [10, 5], we propose a new transformer-based model architecture that includes an image encoder and a tensorial object reconstructor to predict a tensorial SDF field conditioned on multi-view image features
Specifically, we use a vision transformer, DINO2 [1], to encode feature tokens $F_{I}$ from multi-view images as shown in Figure 2. To ensure that the resulting feature tokens incorporate multi-view camera information, we modulate the camera information as described in LRM [5] and inject it into the image encoder using AdaLN [6, 21]. Furthermore, we arrange the tensor vectors $V^{m}_{k}$ and ${M}_{k}^{\tilde{m}}$ into learnable tokens and feed these tokens into a feature decoder composed of a sequence of transformer layers, resulting in a series of tensor tokens. It should be noted that the information from the conditioned image feature $F_{I}$ is connected to the cross-attention layer within these transformer layers. Finally, we reshape and upsample the tensor tokens into the final tensorial SDF representation. Then, all the transformer architecture can be trained in an end-to-end manner using image reconstruction loss at novel views images rendered from the tensorial SDF representation using volume rendering [19]. More details of the network structure can be found in the appendix. The training loss in this stage is defined as,
$\mathcal{L}_{1}$ $=\mathcal{L}_{mse}(I_{rgb},I^{GT}_{rgb})+\mathcal{L}_{mse}(I_{\alpha},I^{GT}_{\alpha})+\mathcal{L}_{mse}(I_{h},I^{GT}_{h})+\lambda_{vgg}\mathcal{L}_{lpips}(I_{rgb},I^{GT}_{rgb})$ (4)
where $I_{rgb},I_{\alpha},I_{\alpha}$ respectively represent the predicted final color image, albedo image, and mask image. Following LRM [5], we apply mean square error loss and VGG-based LPIPS loss [38] to the predicted images and corresponding reference images. During training, we set $\lambda_{\text{vgg}}=2$.
3.4 Lifting SDF to Fine Mesh
Till now, we have introduced a transformer structure that enables the generation of a tensorial SDF representation from sparse images. We can directly extract geometric surfaces from the SDF representation using Marching Cubes [18] after training. However, the strategy has the issues: 1) given the high computational cost of volume rendering, only a patch of the image is rendered during training, thus being unable to fully utilize full-resolution supervisory images. 2) a gradient-based mesh optimization layer after a SDF field can improve the silhouette quality as described in Nvdiffrec [20].
FlexiCubes [26] is a state-of-the-art differentiable iso-surface extraction layer. Employing the differentiable method to extract meshes from the tensorial SDF representation, rasterizing, and rendering at high resolution for supervised learning is beneficial for achieving finer details. One strategy is to use the Flexicube render layer from the outset to replace the computationally expensive volume rendering. However, due to the design of the discrete SDF grid, under sparse view supervision, each gradient descent step affects only the vicinity of the SDF grid vertex. Therefore, directly using Flexicube from the start during training cannot get the network to converge. CRM [33] has similar observations: predicting the correct geometry directly with Flexicube is extremely difficult using pure RGB images. In contrast, volume rendering sampling points throughout the entire space and gradients affect the entire space.
Two-stage training from global to local. Therefore, we combine the advantages of both training with volume rendering and Flexicube and propose a two-stage training strategy. In the first stage, we use the more stable volume rendering for the learning of the global features of the network, allowing the network to achieve global convergence. In the second stage, we initialize the training with the model from the first stage and use Flexicube to optimize the local features, achieving higher resolution textures. During the second stage, to predict a set of weights in each grid cell and the deformation of each grid vertex following Flexicube [26], we additionally introduced two MLP decoders to decode the tensor features from the tensorial SDF:
$d=\Theta_{d}(V_{p},p),w=\Theta_{d}(V_{p},p),$ (5)
where $\Theta_{d},\Theta_{w}$ represent the MLPs used to decode the grid point deformation value and weights of the grid cell, which are needed for the FlexiCube rendering pipeline, respectively. The model is then trained under supervision using the following loss:
$\mathcal{L}_{2}$ $=\mathcal{L}_{1}+\lambda_{d}\left\|I_{d}-I_{d}^{gt}\right\|_{1}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}}$ (6)
where $I_{d}$ represent the rendered depth map, $I_{d}^{gt}$ are the corresponding reference images. The $L_{reg}$ is the regularization term following FlexiCubes [26]. During training, we set $\lambda_{\text{d}}=0.5$, $\lambda_{\text{reg}}=0.005$.
4 Experiment
Figure 4: Qualitative comparison with baselines shows that our method produces high-quality 3D assets with smooth geometry and clear textures, which align well with the input image.
4.1 Implementation Details
Training Datasets
We train the model in GObjaverse from RichDreamer [23], which is rendered using the TIDE renderer on Objaverse [3] and includes G-buffer rendering data such as albedo, RGB, depth, and normal map images from multiple views in resolution of 512$\times$512. Following LGM [30], we utilize a filtered subset, which excludes low-quality 3D models, resulting in a final set of around 80K 3D objects. Specifically, this dataset includes 36 random views of a centered object, with elevations ranging from -5° to 30°, and two additional views for the top and bottom. Although the camera poses for images produced by multi-view diffusion are predetermined, during training, a random subset of 8 images is selected. 4 images are used as inputs to the model, and the remaining 4 are employed for novel view supervision. This strategy not only makes the model robust to inconsistencies in inference inputs but also ensures that the image encoder is sensitive to camera poses. For inference, images with fixed camera poses generated in the first stage are fed to the tensorial SDF reconstruction.
Figure 5: Comparing model training performance across different object representations, Tensorial SDF emerges as the outstanding performer.
4.2 Comparison
We compare our method with the previous state-of-the-art feed-forward generation methods, including One-2-3-45 [14], LRM [5], CRM [33] and LGM [30]. We evaluate our method on the Google Scanned Object (GSO) dataset [4]. To quantitatively evaluate the image quality synthesized by our approach, we adopt three standard metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [34], and Learned Perceptual Image Patch Similarity (LPIPS) [38]. Regarding geometric quality, we report the Chamfer Distance (CD) and Volume Intersection over Union (IoU) as metrics. The quantitative results, as shown in Table 2, demonstrate that our method outperforms others in both color and geometry. Additionally, we show a qualitative comparison of various methods in the Figure 4. As shown in the figure, these methods have all achieved good results, but some of them struggle with certain detailed aspects. One2345 [14] is limited by the discrete SDF volume grid representation, resulting in blurred textures and geometry. LRM [5], lacking multi-view images as constraints, produces 3D models with inconsistent back and front colors and suffers from multi-face Janus problem [28]. LGM [30], using 3DGS as its representation, tends to produce blurred details in areas with dense geometry, such as leaf clusters. CRM [33] exhibits geometric discontinuities in the frog’s hands and the flower’s petioles. However, our method successfully generates 3D assets that are well-aligned with the input conditions, featuring smooth and intact geometric structures.
  Method Tto3d Ito3d Rep. Illum. dec.
LRM $\times$ NeRF $\times$
LGM 3DGS $\times$
CRM $\times$ SDF $\times$
Instand3d NeRF $\times$
Ours TensoSDF
Table 1: The features of different methods, our framework provides comprehensive feature support.
  PSNR ↑ SSIM ↑ LPIPS ↓ CD ↓ IoU ↑
One-2-3-45 18.93 0.779 0.166 0.0614 0.4126
CRM 21.59 0.864 0.159 0.0335 0.4213
OpenLRM 20.29 0.829 0.186 0.0482 0.3731
LGM 20.05 0.798 0.176 0.0417 0.4331
Ours 22.52 0.873 0.143 0.0241 0.4361
Table 2: Quantitative results demonstrate that the color and geometric quality of the 3D assets generated by our method outperforms other methods.
4.3 Ablation Study
Different object representation.
We conducted ablation studies on various object representations to validate their expressiveness and training convergence speed. Note that due to the substantial training costs associated with our final model, the subsequent analytical experiments employ a significantly reduced version of the LDM model. Inspired by previous works [5, 10], we reduced the training dataset to a subset containing 10k objects to accelerate model convergence. We conducted training without image cropping using volume rendering, rendering the images at a fixed resolution of 128$\times$128. Each model was trained on 8 NVIDIA A6000 GPUs over 40 epochs. The quantitative results, as shown in Figure 5, indicate that compared to the representation of Triplane SDF, the Tensorial SDF representation achieves better quality under the same number of epochs of training and converges faster. Furthermore, it can be observed that constructing objects as Tensorial SDF rather than Tensorial NeRF also aids in convergence. We believe this may be because the SDF representation introduces stronger smoothness constraints, reducing geometric disintegration. Additionally, qualitative comparisons of results, as illustrated in Figure 7, demonstrate that Tensorial SDF exhibits superior geometric details with smoother surfaces (e.g., the car hood).
Flexicubes layer
We conducted ablation experiments to verify the impact of introducing the Flexicube layer for model fine-tuning. As shown in  Figure 9 and  Table 7, training with the Flexicubes layer effectively enhances the texture clarity, thanks to the utilization of higher resolution during training. In addition, introducing depth constraints leads to smoother geometric surfaces. For more ablation discussions, please refer to the appendix  Section 6.2.
Figure 6: Qualitative comparison results of Tensorial SDF* and Triplane SDF. Note, for a fair comparison, the Tensorial SDF* used here is a scaled small model aligned with Triplane SDF, not the full model.
  Method PSNR↑ SSIM↑ LPIPS↓
  w/ Flex 23.21 0.869 0.131
w/o Flex 22.76 0.853 0.145
w/ view finetune 22.89 0.856 0.149
w/o view finetune 21.63 0.786 0.157
Figure 7: Quantitative effect of the Flexlayer, evaluate the performance of the fine-tuned model with varying numbers of input views. More discussion can be find in  Section 6.2.
5 Conclusion
This paper introduces a novel feed-forward framework capable of reconstructing 3D meshes with illumination-decoupled textures from a single image input or text prompt in 10 seconds. Our method utilizes a conditional multi-view diffusion model to generate consistent four-view images and lifts them to 3D using a transformer-based large reconstruction model. We propose to generate objects represented as tensorial SDF field, which is more expressive and can accelerate model convergence compared to previous tri-plane representations. In addition, we decompose the rendering color into albedo color and shading color, which enables the generated 3D assets to be easily relit and edited for material properties. We believe that the ability to produce relightable 3D assets is very important for downstream applications.